# MRI motion correction via efficient residual-guided denoising diffusion probabilistic models

### v1

@@ Abstract
Purpose: Motion artifacts in magnetic resonance imaging (MRI) significantly degrade image quality and hinder quantitative downstream applications. Conventional methods to mitigate these artifacts, including repeated acquisitions or motion tracking, impose substantial financial and workflow burdens. This study introduces Res-MoCoDiff, an efficient denoising diffusion probabilistic model specifically designed for MRI motion artifact correction.
Methods: Res-MoCoDiff exploits a novel residual error shifting mechanism during the forward diffusion process to incorporate information from motion-corrupted images. This mechanism allows the model to simulate the evolution of noise with a probability distribution closely matching that of the corrupted data, enabling a reverse diffusion process that requires only four steps. The model employs a U-net backbone, with conventional attention layers replaced by Swin Transformer blocks, to enhance robustness across various image resolutions. Furthermore, the training process integrates a combined $\ell_{1}+\ell_{2}$ loss function, which promotes image sharpness and reduces pixel-level errors. Res-MoCoDiff was evaluated on both an in-silico dataset generated using a realistic motion simulation framework and an in-vivo MR-ART dataset. Comparative analyses were conducted against established methods, including CycleGAN, Pix2pix, and a diffusion model with a vision transformer backbone (MT-DDPM), using quantitative metrics such as peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and normalized mean squared error (NMSE).
Results: The proposed method demonstrated superior performance in removing motion artifacts across minor, moderate, and heavy distortion levels. Res-MoCoDiff consistently achieved the highest SSIM and the lowest NMSE values, with a PSNR of up to $41.91\pm 2.94$ dB for minor distortions. Notably, the average sampling time was reduced to 0.37 seconds per batch of two image slices, compared with 101.74 seconds for conventional approaches.
Conclusion: Res-MoCoDiff offers a robust and efficient solution for correcting MRI motion artifacts, preserving fine structural details while significantly reducing computational overhead. Its rapid processing speed and high restoration fidelity underscore its potential for seamless integration into clinical workflows, ultimately enhancing diagnostic and treatment accuracy and patient care.
@@ 1 Introduction
Magnetic resonance imaging (MRI) is a cornerstone of modern diagnostics, treatment planning, and patient follow-up, providing high-resolution images of soft tissues without the use of ionizing radiation. However, prolonged MRI acquisitions increase the likelihood of patient movement, leading to motion artifacts. These artifacts can alter the B0 field, which results in susceptibility artifacts [1], and disrupt the k-space readout lines, potentially violating the Nyquist criterion and causing ghosting and ringing artifacts [2]. As one of the most common artifacts encountered in MRI [3], motion artifacts may compromise post-processing procedures such as image segmentation [4] and target tracking in MR-guided radiation therapy [5]. Moreover, due to the limited availability of prospective motion correction techniques and the additional complexity they introduce, clinical workflows often resort to repeating the imaging acquisition, which can incur costs exceeding \$300k per scanner annually [6]. The severity and spatial distribution of these artifacts underscore a need for robust methods capable of effectively removing or significantly reducing motion artifacts without repeated imaging.
@@
Traditional motion correction (MoCo) algorithms have primarily aimed at mitigating motion artifacts by optimizing image quality metrics such as entropy and image gradient [7], as well as by estimating the motion-corrupted k-space lines [8] and corresponding motion trajectories [9]. Deep learning (DL) approaches have demonstrated superior performance in removing motion artifacts compared to conventional methods [10, 11]. In particular, both supervised and unsupervised generative DL models based on generative adversarial networks (GANs) have been successfully employed for MRI motion artifact removal [12, 13, 14]. However, GAN-based approaches frequently encounter practical limitations, including mode collapse and unstable training, which can negatively affect the quality of restored images.
@@
Recently, diffusion denoising probabilistic models (DDPMs) have revolutionized image generation techniques by markedly improving synthesis quality [15] and have been adapted for various medical imaging tasks, including image synthesis [16, 17], denoising [18], MRI acceleration [19], and the vision foundation model for 3D MRI [20, 21]. DDPMs involve a forward process in which a Markov chain gradually transforms the input image into Gaussian noise, $\mathcal{N}(\mathbf{0},\mathbf{I})$, over a large number of steps, followed by a reverse process in which a neural network reconstructs the original image from the noisy data [22]. Existing DDPM-based MoCo models concatenate the motion-corrupted image $\mathbf{y}$ with $\mathbf{x}_{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and performing the backward diffusion to reconstruct the motion-free image $\hat{\mathbf{x}}$ being similar to the ground truth image $\mathbf{x}$ [23, 24, 25]. Although these methods achieved promising results, their reliance on numerous diffusion steps substantially increases the inference time. Additionally, initiating reconstruction from fully Gaussian noise $\mathbf{x}_{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ might be suboptimal for MRI motion correction task (see Section 2.1).
@@
In this study, we present Res-MoCoDiff, a diffusion model that explicitly exploits the residual error between motion-free $\mathbf{x}$ and motion-corrupted $\mathbf{y}$ images (i.e., $\mathbf{r}=\mathbf{y}-\mathbf{x}$) in the forward diffusion process. Integrating this residual error into the diffusion process enables generation of noisy images at step $N$ with a probability distribution closely matching that of the motion-corrupted images, specifically $p(\mathbf{x}_{N})\sim\mathcal{N}(\mathbf{x};\mathbf{y},\gamma^{2}\mathbf{I})$. This approach offers two significant advantages: (1) enhanced reconstruction fidelity by avoiding the restrictive purely Gaussian prior assumption of conventional DDPMs, and (2) substantial computational efficiency, as the reverse diffusion process can be reduced to only four steps, substantially accelerating reconstruction times compared to traditional DDPMs. Throughout this paper, we refer to our method as the residual-guided efficient motion-correction denoising diffusion probabilistic model, or “Res-MoCoDiff”.
@@ 2.1 DDPM
DDPMs are inspired by non-equilibrium thermodynamics and aim to approximate complex data distributions using a tractable distribution, such as a normal Gaussian distribution as a prior [26]. Specifically, DDPMs employ a Markov chain consisting of two distinct processes: a forward diffusion and a backward (denoising) process. During the forward diffusion, the input image $\mathbf{x}$ is gradually perturbed through a sequence of small Gaussian noise injections, eventually converging toward pure Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$ after a large number of diffusion steps [22, 15]. Conversely, the backward process employs a deep learning (DL) model to iteratively remove noise and reconstruct the original image from the Gaussian noise by approximating the reverse Markov chain of the forward diffusion.
@@ 2.2.2 Backward process
[…]
@@
motion-corrupted image $\mathbf{y}$ $\mathbf{x}_{N}\sim\mathcal{N}(\mathbf{x}_{N};\mathbf{y},\gamma^{2}\beta_{N}\mathbf{I})$
for t = N, …, 1 do $\bm{\epsilon}\sim\mathcal{N}(\bm{\epsilon};\mathbf{0},\mathbf{I})$ if $t>1$ else $\bm{\epsilon}=0$ $\bm{\mu}_{\theta}=\dfrac{\beta_{t-1}}{\beta_{t}}\mathbf{x}_{t}+\dfrac{\alpha_{t}}{\beta_{t}}f_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$triangleright Given in (10) $\mathbf{x}_{t-1}=\bm{\mu}_{\theta}+\gamma\sqrt{\dfrac{\beta_{t-1}\alpha_{t}}{\beta_{t}}}\bm{\epsilon}$
@@
Res-MoCoDiff was implemented in PyTorch (version 2.5.1) and executed on an NVIDIA A100 GPU. The model was trained for 100 epochs with a batch size of 32. Optimization was performed using the RAdam optimizer [32] in conjunction with a cosine annealing learning rate scheduler [33]. A warm-up phase comprising 5,000 steps was employed prior to transitioning to the cosine schedule to stabilize early training dynamics.
@@ 2.3 Patient Data Acquisition and Data Pre-processing
The IXI dataset comprises 580 cases of T1-weighted (T1-w) brain MRI images. We partitioned the dataset into two non-overlapping subsets: a training set consisting of 480 cases (54,160 slices) and a testing set comprising 100 cases (11,980 slices). We adapted the motion simulation technique of Duffy et al. [11] to generate an in-silico dataset with varying levels of motion artifacts including high, moderate, and minor by perturbing 15, 10, and 7 k-space lines along a phase encoding direction, respectively. Random slabs, with widths ranging between three and seven k-space lines, were selected along the phase encoding direction and were subjected to rotational perturbations of ±$\pm 7^{\circ}$ and translational shifts of ±$\pm 5$ mm.
Additionally, model performance on in-vivo data was evaluated using the MR-ART dataset, which comprises 148 cases (95 females and 53 males). This dataset includes three types of images: ground truth motion-free images, motion-corrupted images with a low level of distortion (level 1), and motion-corrupted images with a high level of distortion (level 2). Rigid brain image registration was performed using FSL-FLIRT [35, 36] to compensate for misalignment between the motion-free and motion-corrupted images.
@@ 2.4 Quantitative and Statistical Analysis
To quantitatively assess the performance of the models in removing brain motion artifacts, we reported three metrics: normalized mean squared error (NMSE), structural similarity index measure (SSIM) [40], and peak signal-to-noise ratio (PSNR). Lower NMSE values indicate better performance, although NMSE may favor solutions with increased blurriness [41]. SSIM ranges from -1 to 1, with a value of 1 representing optimal structural similarity between the reconstructed and ground truth images. Likewise, a higher PSNR denotes improved performance and is more aligned with human perception due to its logarithmic scaling [42]. The quantitative metrics were computed using the PIQ library (version 0.8.0) [43] with its default parameters.
@@ 3.1 Qualitative results
The motion artifacts observed in the motion-corrupted images confirm that our simulation procedure successfully reproduces both ringing artifacts inside the skull and ghosting of bright fat tissue outside the skull, as indicated by the white and green arrows in Figure 4(a) and (d). Notably, the zoomed-in regions in Figure 4(b) illustrate that Res-MoCoDiff preserves fine structural details more effectively than the comparative methods. Furthermore, the pixel-level distortion maps in Figure 4(c), (f), and (i) underscore the superior artifact removal achieved by Res-MoCoDiff.
@@
Although our approach demonstrates a generally robust ability to preserve detailed structures, a few residual ringing artifacts remain (highlighted by arrows in Figure 4(d)) for the moderate distortion level. For the minor distortion level, the overall performance among all methods is similar in mitigating motion artifacts, as shown in Figure 4(g)-(i). Finally, the pixel-wise correlation plots in Figure 4(j)-(l) confirm the qualitative findings: Res-MoCoDiff attains Pearson correlation coefficients of $\rho=0.9974$, $\rho=0.9990$, and $\rho=0.9999$ for high, moderate, and minor distortion levels, respectively, surpassing the second-best MT-DDPM method, which yields $\rho=0.9961$, $\rho=0.9987$, and $\rho=0.9997$.
@@
Figure 4: Qualitative results for the in-silico dataset are shown. Panels (a)-(c), (d)-(f), and (g)-(i) illustrate the outcomes for heavy, moderate, and minor distortion levels, respectively. Panels (j)-(l) display the corresponding pixel-wise correlation plots for heavy, moderate, and minor distortion levels. The white and green arrows in panels (a) and (d) indicate ringing artifacts inside the skull and ghosting of bright fat tissue outside the skull, respectively, while panels (b), (e), and (h) present zoomed-in views of the regions highlighted by the red boxes.
@@
Figure 5: The qualitative results for the in-vivo dataset are illustrated. The green and red arrows indicate the ringing artifacts inside the skull, and the white arrow shows the ghosting of bright fat tissue outside of the skull.
Qualitative results for the in-vivo dataset are presented in Figure 5, where the top row corresponds to Level 1 motion corruption and the bottom row to Level 2. In both levels, red arrows mark ringing artifacts within the skull, and a white arrow highlights ghosting of bright fat tissue outside the skull (particularly visible at Level 2). The green arrows denote regions where Res-MoCoDiff successfully recovers fine structural details. Accompanying PSNR and SSIM values—improved from $28.81$ dB and 75.71 (Defective) to $30.61$ dB and 95.30 (Res-MoCoDiff) for Level 1, and from $27.03$ dB and 67.58 (Defective) to $30.78$ dB and 94.75 (Res-MoCoDiff) for Level 2, which further validate the enhanced performance achieved by our proposed method.

### v2

# Res-MoCoDiff: Residual-guided diffusion models for motion artifact correction in brain MRI

@@ Abstract
Objective.Motion artifacts in brain MRI, mainly from rigid head motion, degrade image quality and hinder downstream applications. Conventional methods to mitigate these artifacts, including repeated acquisitions or motion tracking, impose workflow burdens. This study introduces Res-MoCoDiff, an efficient denoising diffusion probabilistic model specifically designed for MRI motion artifact correction.Approach.Res-MoCoDiff exploits a novel residual error shifting mechanism during the forward diffusion process to incorporate information from motion-corrupted images. This mechanism allows the model to simulate the evolution of noise with a probability distribution closely matching that of the corrupted data, enabling a reverse diffusion process that requires only four steps. The model employs a U-net backbone, with attention layers replaced by Swin Transformer blocks, to enhance robustness across resolutions. Furthermore, the training process integrates a combined $\ell_{1}+\ell_{2}$ loss function, which promotes image sharpness and reduces pixel-level errors. Res-MoCoDiff was evaluated on both an in-silico dataset generated using a realistic motion simulation framework and an in-vivo MR-ART dataset. Comparative analyses were conducted against established methods, including CycleGAN, Pix2pix, and a diffusion model with a vision transformer backbone (MT-DDPM), using quantitative metrics such as peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and normalized mean squared error (NMSE).Main results. The proposed method demonstrated superior performance in removing motion artifacts across minor, moderate, and heavy distortion levels. Res-MoCoDiff consistently achieved the highest SSIM and the lowest NMSE values, with a PSNR of up to $41.91\pm 2.94$ dB for minor distortions. Notably, the average sampling time was reduced to 0.37 seconds per batch of two image slices, compared with 101.74 seconds for conventional approaches.Significance.Res-MoCoDiff offers a robust and efficient solution for correcting MRI motion artifacts, preserving fine structural details while significantly reducing computational overhead. Its speed and restoration fidelity underscore its potential for integration into clinical workflows, enhancing diagnostic accuracy and patient care.
@@ 1 Introduction
Magnetic resonance imaging (MRI) is a cornerstone of modern diagnostics, treatment planning, and patient follow-up, providing high-resolution images of soft tissues without the use of ionizing radiation. However, prolonged MRI acquisitions increase the likelihood of patient movement, leading to motion artifacts. These artifacts can alter the B0 field, which results in susceptibility artifacts [1], and disrupt the k-space readout lines, potentially violating the Nyquist criterion and causing ghosting and ringing artifacts [2]. As one of the most common artifacts encountered in MRI [3], motion artifacts may compromise post-processing procedures such as image segmentation [4] and target tracking in MR-guided radiation therapy [5]. Moreover, due to the limited availability of prospective motion correction techniques and the additional complexity they introduce, clinical workflows often resort to repeating the imaging acquisition. The severity and spatial distribution of these artifacts underscore a need for robust methods capable of effectively removing or significantly reducing motion artifacts without repeated imaging.
Traditional motion correction (MoCo) algorithms have primarily aimed at mitigating artifacts by optimizing image quality metrics such as entropy and image gradient [6], as well as by estimating the motion-corrupted k-space lines [7] and corresponding motion trajectories [8]. Additional strategies include navigator echoes, external tracking devices, and sequence modifications that prospectively adjust acquisition to account for motion [2]. Retrospective approaches in the reconstruction domain, such as phase shift correction in k-space, blind motion trajectory estimation, and compressed sensing with motion models, have also been widely explored [6]. While these methods have shown clinical utility, their adoption is often limited by the need for raw k-space data, which is not routinely stored in clinical archives, and by reconstruction pipelines that vary across scanner vendors, reducing reproducibility and generalizability [9, 10].
@@
In parallel, deep learning (DL) approaches have demonstrated superior performance in suppressing motion artifacts compared to conventional methods [11, 12]. In particular, both supervised and unsupervised generative models based on generative adversarial networks (GANs) have been successfully employed for MRI motion artifact removal [13, 14, 15]. However, GAN-based approaches frequently encounter practical limitations, including mode collapse and unstable training, while reconstruction- and k-space-based methods often require scanner-specific modifications. These limitations highlight the importance of image-domain methods that operate directly on magnitude images, which are universally available and readily integrated into existing clinical workflows without changes to acquisition hardware or vendor-specific reconstruction software.
In contrast, image-domain methods that operate directly on reconstructed magnitude images that are universally available in both retrospective studies and multi-center collaborations, enabling wide applicability without requiring changes to the acquisition hardware or reconstruction software. This characteristic makes image-based models particularly attractive as off-the-shelf solutions for clinical deployment. Furthermore, the use of image-domain methods allows integration with existing clinical pipelines for tasks such as segmentation, registration, and radiotherapy planning, where corrected magnitude images can be seamlessly substituted for corrupted ones. For these reasons, this study focused on the image domain, aiming to balance methodological innovation with clinical feasibility.
@@
Recently, diffusion denoising probabilistic models (DDPMs) have revolutionized image generation techniques by markedly improving synthesis quality [16] and have been adapted for various medical imaging tasks, including image synthesis [17, 18], denoising [19], MRI acceleration [20], and the vision foundation model for MRI [21, 22, 23]. DDPMs involve a forward process in which a Markov chain gradually transforms the input image into Gaussian noise, $\mathcal{N}(\mathbf{0},\mathbf{I})$, over a large number of steps, followed by a reverse process in which a neural network reconstructs the original image from the noisy data [24]. Existing DDPM-based MoCo models concatenate the motion-corrupted image $\mathbf{y}$ with $\mathbf{x}_{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and performing the backward diffusion to reconstruct the motion-free image $\hat{\mathbf{x}}$ being similar to the ground truth image $\mathbf{x}$ [25, 26, 27]. Although these methods achieved promising results, their reliance on numerous diffusion steps substantially increases the inference time. Additionally, initiating reconstruction from fully Gaussian noise $\mathbf{x}_{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ might be suboptimal for MRI motion correction task (see Section 2.1).
@@
In this study, we present Res-MoCoDiff, residual-guided efficient motion-correction denoising diffusion probabilistic model, a diffusion model that explicitly exploits the residual error between motion-free $\mathbf{x}$ and motion-corrupted $\mathbf{y}$ images (i.e., $\mathbf{r}=\mathbf{y}-\mathbf{x}$) in the forward diffusion process. Integrating this residual error into the diffusion process enables generation of noisy images at step $N$ with a probability distribution closely matching that of the motion-corrupted images, specifically $p(\mathbf{x}_{N})\sim\mathcal{N}(\mathbf{x};\mathbf{y},\gamma^{2}\mathbf{I})$. This approach offers two significant advantages: (1) enhanced reconstruction fidelity by avoiding the restrictive purely Gaussian prior assumption of conventional DDPMs, and (2) substantial computational efficiency, as the reverse diffusion process can be reduced to only four steps, substantially accelerating reconstruction times compared to traditional DDPMs.
@@ 2.1 DDPM
DDPMs are inspired by non-equilibrium thermodynamics and aim to approximate complex data distributions using a tractable distribution, such as a normal Gaussian distribution as a prior [28]. Specifically, DDPMs employ a Markov chain consisting of two distinct processes: a forward diffusion and a backward (denoising) process. During the forward diffusion, the input image $\mathbf{x}$ is gradually perturbed through a sequence of small Gaussian noise injections, eventually converging toward pure Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$ after a large number of diffusion steps [24, 16]. Conversely, the backward process employs a DL model to iteratively remove noise and reconstruct the original image from the Gaussian noise by approximating the reverse Markov chain of the forward diffusion.
@@ 2.2.2 Backward process
$\centering\hat{\theta}=\underset{\theta}{\arg\min}\parallel f_{\theta}(\mathbf{x}_{t},\mathbf{y},t)-\mathbf{x}\parallel_{2}^{2},\@add@centering$ (11)
@@
In our implementation, the training objective combined $\ell_{1}$ and $\ell_{2}$ losses with equal weighting. This design was chosen to balance the strengths of the two norms: the $\ell_{2}$ component penalizes large deviations and promotes overall pixel-level consistency, whereas the $\ell_{1}$ component preserves high-frequency features and sharper structural details. While $\ell_{2}$ loss alone can lead to overly smoothed reconstructions, the inclusion of $\ell_{1}$ mitigates this effect by encouraging sparsity and edge sharpness. This complementary behavior has been reported in prior medical image restoration studies and was confirmed in our ablation analysis (Section 3.3). No additional hyperparameter optimization of the relative weights was performed in this work.
In Res-MoCoDiff, the forward diffusion process was implemented with $N=20$ steps. However, the reverse process was reduced to only four steps due to the residual-guided formulation, which shifts the corrupted image distribution closer to the motion-free distribution. This substantially reduces the gap that the reverse diffusion must bridge. The number of reverse steps was selected empirically by evaluating different step counts on a validation set, where four steps provided the optimal trade-off between reconstruction quality and efficiency.
@@
motion-corrupted image $\mathbf{y}$; number of steps $N=4$; noise scheduler $\{\beta_{t}\}_{t=1}^{N}$ with $\beta_{N}=0.999$, $\beta_{1}=(0.04/\gamma)^{2}$, growth rate $p=0.3$, and $\alpha_{t}=\beta_{t}-\beta_{t-1}$; noise scaling $\gamma=2$ $\mathbf{x}_{N}\sim\mathcal{N}(\mathbf{x}_{N};\mathbf{y},\gamma^{2}\beta_{N}\mathbf{I})$
for t = N, …, 1 do $\bm{\epsilon}\sim\mathcal{N}(\bm{\epsilon};\mathbf{0},\mathbf{I})$ if $t>1$ else $\bm{\epsilon}=0$ $\bm{\mu}_{\theta}=\dfrac{\beta_{t-1}}{\beta_{t}}\mathbf{x}_{t}+\dfrac{\alpha_{t}}{\beta_{t}}f_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$ $\triangleright$ Given in (10) $\mathbf{x}_{t-1}=\bm{\mu}_{\theta}+\gamma\sqrt{\dfrac{\beta_{t-1}\alpha_{t}}{\beta_{t}}}\bm{\epsilon}$
@@
Res-MoCoDiff was implemented in PyTorch (version 2.5.1) and executed on an NVIDIA A100 with 80 GB GPU RAM. The model was trained for 100 epochs with a batch size of 32. Optimization was performed using the RAdam optimizer [34] in conjunction with a cosine annealing learning rate scheduler [35], with an initial learning rate of $2\times 10^{-4}$ and a minimum learning rate of $2\times 10^{-5}$. A warm-up phase comprising 5,000 steps was employed prior to transitioning to the cosine schedule to stabilize early training dynamics.
@@ 2.3 Patient Data Acquisition and Data Pre-processing
The IXI dataset comprises 580 cases of T1-weighted (T1-w) brain MRI images. We partitioned the dataset into two non-overlapping subsets: a training set consisting of 480 cases (54,160 slices) and a testing set comprising 100 cases (11,980 slices). We adapted the motion simulation technique of Duffy et al. [12] to generate an in-silico dataset with varying levels of motion artifacts including high, moderate, and minor by perturbing 15, 10, and 7 k-space lines along a phase encoding direction, respectively. Random slabs, with widths ranging between three and seven k-space lines, were selected along the phase encoding direction and were subjected to rotational perturbations of $\pm 7^{\circ}$ and translational shifts of $\pm 5$ $m$.
Additionally, model performance on in-vivo data was evaluated using the MR-ART T1-w brain MRI dataset, which comprises 148 cases (95 females and 53 males). This dataset includes three types of images: ground truth motion-free images, motion-corrupted images with a low level of distortion (level 1), and motion-corrupted images with a high level of distortion (level 2). Rigid brain image registration was performed using FSL-FLIRT [37, 38] to compensate for misalignment between the motion-free and motion-corrupted images.
It is important to note that our study was conducted in the image domain, rather than in k-space. This choice was made because most publicly available datasets and many clinical archives provide only reconstructed magnitude images, not raw k-space data. Working in the image domain therefore enables broader applicability of Res-MoCoDiff to retrospective studies and multi-institutional data, where access to raw acquisition information is not readily available.
@@ 2.4 Quantitative and Statistical Analysis
To quantitatively assess the performance of the models in removing brain motion artifacts, we reported three metrics: normalized mean squared error (NMSE), structural similarity index measure (SSIM) [42], and peak signal-to-noise ratio (PSNR). Lower NMSE values indicate better performance, although NMSE may favor solutions with increased blurriness [43]. SSIM ranges from -1 to 1, with a value of 1 representing optimal structural similarity between the reconstructed and ground truth images. Likewise, a higher PSNR denotes improved performance and is more aligned with human perception due to its logarithmic scaling [44]. The quantitative metrics were computed using the PIQ library (https://piq.readthedocs.io/en/latest, version 0.8.0) [45] with its default parameters.
@@ 3.1 Qualitative results
The motion artifacts observed in the motion-corrupted images confirm that our simulation procedure successfully reproduces both ringing artifacts inside the skull and ghosting of bright fat tissue outside the skull, as indicated by the white and green arrows in Figures 4(a) and (d). Notably, the zoomed-in regions in Figure 4(b) illustrate that Res-MoCoDiff preserves fine structural details more effectively than the comparative methods. Furthermore, the pixel-level distortion maps in Figures 4(c), (f), and (i) underscore the superior artifact removal achieved by Res-MoCoDiff.
@@
Although our approach demonstrates a generally robust ability to preserve detailed structures, a few residual ringing artifacts remain (highlighted by arrows in Figure 4(d)) for the moderate distortion level. For the minor distortion level, the overall performance among all methods is similar in mitigating motion artifacts, as shown in Figure 4(g)-(i). Finally, the pixel-wise correlation plots in Figure 5(a)-(c) confirm the qualitative findings: Res-MoCoDiff attained Pearson correlation coefficients of $\rho=0.9974$, $\rho=0.9990$, and $\rho=0.9999$ for high, moderate, and minor distortion levels, respectively, surpassing the second-best MT-DDPM method, which yields $\rho=0.9961$, $\rho=0.9987$, and $\rho=0.9997$.
@@
Figure 4: Qualitative results for the in-silico dataset are shown. Panels (a)-(c), (d)-(f), and (g)-(i) illustrate the outcomes for heavy, moderate, and minor distortion levels, respectively. The white and green arrows in panels (a) and (d) indicate ringing artifacts inside the skull and ghosting of bright fat tissue outside the skull, respectively, while panels (b), (e), and (h) present zoomed-in views of the regions highlighted by the red boxes.
Figure 5: Pixel-wise correlations for the in-silico dataset are shown. Panels (a)-(c) display the corresponding pixel-wise correlation plots for heavy, moderate, and minor distortion levels.
@@
Figure 6: Qualitative results for the in-vivo MR-ART dataset, illustrating both optimal (a) and suboptimal (b) reconstructions. Green and red arrows indicate ringing artifacts inside the skull, while the white arrow highlights ghosting of bright fat tissue outside the skull. Yellow arrows denote regions where Res-MoCoDiff hallucinated structures due to insufficient soft-tissue contrast in the corrupted inputs.
Qualitative results for the in-vivo MR-ART dataset are presented in Figure 6, where both optimal and suboptimal reconstructions are illustrated. In panel (a), corresponding to cases with recoverable motion corruption, red arrows indicate ringing artifacts inside the skull, and a white arrow highlights ghosting of bright fat tissue outside the skull. The green arrows denote regions where Res-MoCoDiff successfully restores fine structural details, achieving improvements from 28.81 dB and 75.71 (motion-corrupted) to 30.61 dB and 95.30 (Res-MoCoDiff) in PSNR and SSIM for Level 1, and from 27.03 dB and 67.58 to 30.78 dB and 94.75 for Level 2. In panel (b), however, we show suboptimal reconstructions in which Res-MoCoDiff failed to fully recover the anatomical structures. As indicated by the yellow arrows, the model produced hallucinated features when the motion-corrupted input lacked sufficient soft-tissue contrast, underscoring the inherent difficulty of reconstructing details that are not present in the corrupted images.